How many years will the following take?
$210 to grow to $21,908 if invested at 4.79 percent, compounded annually
Answer #1

Using Future value formula Calculated number of periods

FV= future value = 21908

PV= present value = 210

r= rate of return = 4.79

n= number of periods

FV= PV (1+r)^n

21908= 210 (1+0.0479) ^n

210 (1+0.0479) ^n= 21908 / 210

In (1+0.0479) ^n = In 104.323809

n= 4.647499610 /0.046788161

n= 99.331

Note: both side used natural logarithm (In) and n= 99.331 is correct answer to get 21908 future value or Present Value 210